# Destructive and constructive RIS beamforming in an ISAC-multi-user MIMO network

### v1

@@ IV Malicious RIS sensing SNR degradation
The simulation results are averaged over $500$ independent channel realizations with $\left|\mathcal{Z}\right|=\left|\mathcal{R}\right|=1000$. The number of BS antennas is $M=30$ and the number of UEs is $K=4$ and, unless otherwise specified, $N=15$, with a number of faulty pixels equal to $\left|\mathcal{Q}\right|=4$. The static paths $\mathbf{h}_{s,k}$, $\mathbf{h}_{s,{\rm T}}$, and $\mathbf{g}$ are independent and identically distributed (i.i.d.) Rician fading with a K-factor of $10$ whereas $\mathbf{H}_{t}$ is a scattering-free line of sight (LoS) channel. Lastly, $\mathbf{h}_{r,k}$ are i.i.d Rayleigh fading. The static and reflected path RCS variances are set to $\delta_{r}^{2}=\delta_{s}^{2}=10^{-5}$ whereas $\delta_{m}=9\cdot 10^{-6}$. We assume an equal minimum SINR for all UEs, denoted by $\gamma$ and, unless otherwise specified, it is equal to $2$. The BS available power is $3$ dB and the UEs and sensing noise variances are set to $\sigma_{k}^{2}=\sigma_{\rm T}^{2}=1$. We consider the $2$D area shown in Fig. 2, where the UEs are randomly distributed within the highlighted area.
Refer to caption(0,0)00(0,0)( 0 , 0 )$20$m20202020m60606060m20202020my50505050m50505050m50505050mxBSRISUEsTarget
@@ V-A Sensing SNR maximization
We have first tested the sensing SNR maximization capabilities of Algorithm 1. Fig. 3 shows that $\rho$ is a monotonically increasing function of the number of iterations $t$ and convergence is reached in $20$ iterations. Unsurprisingly, a larger RIS size allows us to reflect more energy towards the target, resulting in a higher $\rho$. There are diminishing returns in increasing $N$: when $N$ goes from $15$ to $20$, the final value of $\rho$ increases by $5$ dB whereas when $N$ goes from $20$ to $25$ the increase is halved. Fig. 3 shows the impact of clustered pixel failures [8] onto Algorithm 1: We notice that this kind of failure does not alter the maximization trend, but leads to convergence to a lower $\rho$-value. For example, a big phase bias magnitude of $\pi/3$ attains a $\rho$ value only $2$ dB lower than the one obtained with a perfectly functioning RIS, giving our algorithm a certain degree of resilience.
@@ V-B Malicious RIS action
Fig. 5 shows how Algorithm 2 causes SNR reduction. One iteration is sufficient to obtain a $40$ dB SNR reduction. This reduction further increases in the span of 15 iterations, reaching a final value around $-22$ dB. Such an attack is particularly complicated to detect as the communication SINRs are untouched, thus the communication UEs will not see any difference in their service. A reduction of $\gamma$ would improve the SNR reduction at the expenses of a higher vulnerability to detection. However, a $50\%$ reduction in $\gamma$ corresponds to a mere $2$ dB SNR reduction. Furthermore, setting $\gamma$ to $0.5$ gives rise to an SNR that’s $1$dB lower than the one obtained with an $\gamma=1$. We attribute the bump at iteration 6 of the SNR curve with $\gamma=0,5$ to an outlying realization amongst the averaged ones: it is worth mentioning that said bumb does not hinder the SNR convergence process and $\rho$ at iteration $15$ is lower than $\rho$ at iteration $5$. Fig. 6 assesses the impact of clustered RIS pixel failures onto Algorithm 2’s $\rho$ minimization action. Contrary to what we observed in Fig. 4, faulty RIS hardware has a bigger impact onto the final sensing SNR. A phase-shift bias of $\pi/6$ is sufficient to induce a $10$ dB penalty onto the final value of $\rho$, making this algorithm much less resilient than the previous one.
@@ VI Conclusions
We have investigated the potential of a RIS in an ISAC MU-MIMO network. The positive side of the RIS’s presence has been investigated by devising an optimization algorithm maximizing the sensing target SNR under the communication SINR constraints and a finite power budget. The negative side arises when a hacker takes control of the RIS, computing its phase-shifts to minimize the target SNR, under the same SINR constraints. The hacker can make the sensing target practically invisible to the system. We have further taken into account the presence of clustered pixel failures affecting the RIS. Numerical simulations have exemplified how great both potentials are and how the RIS element impairments impact them based on the RIS intentions.

### v3

@@ V Numerical Results
In this section, we will demonstrate the ability of the RIS to either improve or degrade the system performance. The simulation results are averaged over $500$ independent channel realizations with $\left|\mathcal{Z}\right|=\left|\mathcal{R}\right|=1000$. The number of BS antennas is $M=30$ and the number of UEs is $K=4$. Unless otherwise specified, $N=15$, with a number of faulty pixels equal to $\left|\mathcal{Q}\right|=4$. The static paths $\mathbf{h}_{\textrm{s},k}$, $\mathbf{h}_{\textrm{s,T}}$, and $\mathbf{g}$ are independent and identically distributed (i.i.d.) Rician fading with a K-factor of $10$, whereas $\mathbf{H}_{t}$ is a scattering-free line of sight (LoS) channel. Lastly, $\mathbf{h}_{\textrm{r},k}$ are i.i.d. Rayleigh fading. The static and reflected path RCS variances are set to $\delta_{r}^{2}=\delta_{s}^{2}=10^{-5}$ whereas $\delta_{m}=9\cdot 10^{-6}$. We assume an equal minimum SINR for all UEs, denoted by $\gamma$ and it is equal to $2$, unless otherwise specified. The BS available power is $3$ dBW and the UEs and sensing noise variances are set to $\sigma_{k}^{2}=\sigma_{\rm T}^{2}=1$. We consider the $2$D area shown in Fig. 2, where the UEs are randomly distributed within the highlighted area.
@@ V-A Sensing SNR maximization
We will first test the sensing SNR maximization capabilities of Algorithm 1. Fig. 3 shows that $\rho$ is a monotonically increasing function of the number of iterations $t$ and convergence is reached in $20$ iterations. A larger RIS size allows us to reflect more energy towards the target, resulting in a higher $\rho$. There are diminishing returns in increasing $N$. When $N$ goes from $15$ to $20$, the final value of $\rho$ increases by $5$ dB, whereas when $N$ goes from $20$ to $25$ the increase is halved. Fig. 3 shows the impact of clustered-biased pixel failures [8] onto Algorithm 1. This kind of failure does not alter the maximization trend, but leads to convergence to a lower $\rho$-value. For example, a big phase bias magnitude of $\pi/3$ attains a $\rho$ value only $2$ dB lower than the one obtained with a perfectly functioning RIS, showing that nonideal hardware can be tolerated by the system.
@@ V-B Malicious RIS action
Fig. 5 shows how Algorithm 2 causes SNR reduction. One iteration is sufficient to obtain a $40$ dB SNR reduction w.r.t. to the maximized value obtained in the previous section. This reduction further increases in the span of 15 iterations, reaching a final value around $-22$ dB. Such an attack is complicated to detect as the communication SINRs are untouched, thus the communication UEs will not see any difference in their service. Reducing $\gamma$ would improve the SNR reduction at the expense of a higher chance of the attack being discovered. However, a $50\%$ reduction in $\gamma$ corresponds to a mere $2$ dB SNR reduction. We attribute the bump at iteration 6 of the SNR curve with $\gamma=0,5$ to an outlying realization among the averaged ones: it is worth mentioning that said bump does not hinder the SNR convergence process and $\rho$ at iteration $15$ is lower than $\rho$ at iteration $6$. Fig. 6 assesses the impact of the considered pixel failures model onto Algorithm 2. Contrary to what we observed in Fig. 4, faulty RIS hardware has a bigger impact on the final sensing SNR. A bias of $\pi/6$ is sufficient to induce a $10$ dB penalty onto the final value of $\rho$, making this algorithm less tolerant towards faulty hardware.
@@ VI Conclusions
We have investigated the potential benefits and harm of having a RIS in an ISAC MU-MIMO network. The positive side of the RIS’s presence has been investigated by devising an optimization algorithm maximizing the sensing target SNR under the communication SINR constraints and a finite power budget. The negative side arises when a hacker takes control of the RIS, computing its phase-shifts to minimize the target SNR, under the same SINR constraints. The hacker can make the sensing target practically invisible to the system, which is a major concern if ISAC is meant to be used for surveillance. We have further considered the presence of clustered pixel failures in the RIS. Numerical simulations have exemplified how great both potentials are and how the investigated pixel failures caused the final SNR value to be at a fixed offset from the one obtained with an unimpaired RIS.